which is an inequality
x^2/4^2-y^2/5^2 < or eqaul to 1
a. (-8,-2)
b. (2,0)
c. (7,1)
d. (-14,3)

Answer by stanbon(75887)

which point satisfies the inequality?
--------------------------------------------
[x^2/4^2] - [y^2/5^2] <= 1
Test each point:
------
a. (-8,-2)
64/16 - 4/25 <= 1
4 - 4/25 <= 1
100 - 4 <= -25
False
------------------------
b. (2,0)
[4/16] - [0] <= 1
False
--------------------------
c. (7,1)
[49/16] - [1/25] <= 1
False
---------------------
d. (-14,3)
(14^2/16] - [9/25] <= 1
12.25 - (9/25) <= 1
False
===================
Cheers,
Stan H.